Assets 2012 2013
Cash 45 50
A/R 362 301
Inventory 386 237
Total 793 588
Net Fixed Assets 1000 1200
Total Assets 1793 1788
Liabilities 2012 2013
Accounts Payable 341 327
Notes Payable 150 125
Total 491 452
Long-Term Debt 500 750
Common Stock 400 400
Retained Earnings 95 125
1486 1727
2013
Sales 1553
Cost of Goods Sold 1242
Depreciation 200
EBIT 111
Interest 80
Taxes 60
Net Income -29

A. What is the Cash Cycle (to 1 decimal)?

B. What is the Net Working Capital (to nearest dollar with no decimals)?
Answer #1
Answer A
Cash cycle is an efficiency ratio which measures the number of days for which a company’s cash is tied up in inventories and accounts receivable less
days for which the amount payable against those raw materials remained outstanding.
The formula to calculate Cash Cycle is as under,
Cash Cycle = Days Sales Outstanding + Days Inventories Outstanding - Days Payable Outstanding
days sales outstanding = Average Accounts Receivable × 365 ÷ Credit Sales
Average accounts receivables = [$362 + $301]/2 = $331.50
days sales outstanding = $331.50 × 365 ÷ $1553 = 77.9 days
days inventory outstanding = Average Inventories × 365 ÷ Cost of Goods Sold
Average Inventories = [$386 + $237]/2 = $311.50
days inventory outstanding = $311.50 × 365 ÷ $1242 = 91.5 days
days payables outstanding = Average Accounts Payable × 365 ÷ Cost of Goods Sold
Average Accounts Payable = [$341 + $327]/2 = $334
days payables outstanding = $334 × 365 ÷ $1242 = 98.2 days
Cash Cycle = $77.9 days + $91.5 days - $98.2 days = 71.3 days
Answer B
The formula to calculate net working capital is as under,
Net Working capital = Current assets 2013- Current liabilities 2013
Net Working capital = $588 - $452 = $136